OpenStudy (anonymous):

Solve 2 log7 5 + log7 x = log7 100

Nnesha (nnesha):

log rules \[\large\rm \log_b x + \log_b = lob_b ( x \cdot y)\] addition to multiplication ^ \[\large\rm \log_b x - \log_b = lob_b ( \frac{x}{ y})\] subtraction <----> division ^ \[\large\rm \log_b x = lob_b y ~~~~~ \rightarrow \cancel{\log_b} x = \cancel{\log_b} y~~~ \rightarrow x=y\]

OpenStudy (anonymous):

y missed there above from second terms

Nnesha (nnesha):

so which rule would you apply ?

Nnesha (nnesha):

oh yeah thanks

OpenStudy (anonymous):

np

OpenStudy (anonymous):

Wait so would i use the addition to multiplication rule? I'm sorry im terrible at this stuff

Nnesha (nnesha):

log rules \[\large\rm \log_b \color{Red}{x} + \log_b \color{blue}{y} = log_b ( \color{ReD}{x} \cdot \color{blue}{y})\] addition to multiplication ^ \[\large\rm \log_b \color{reD}{x} - \log_b \color{blue}{ y}= log_b ( \frac{\color{red}{x}}{\color{blue}{ y}})\] subtraction <----> division ^ \[\large\rm \color{pink}{\log_b} \color{Red}{x} = \color{pink}{log_b} \color{blue}{y} ~~~~~ \rightarrow \cancel{\color{pink}{\log_b}} \color{red}{ x} = \cancel{\color{pink}{\log_b}}\color{blue}{ y}~~~ \rightarrow x=y\]

Nnesha (nnesha):

yes that's correct

Nnesha (nnesha):

sorry about these typos...

Nnesha (nnesha):

\[ \log_7 5 + \log_7 x = \log_7 100\] is this your question ?

Nnesha (nnesha):

or is it \[ 2\log_7 5 + \log_7 x = \log_7 100\] ?

OpenStudy (anonymous):

So would it be something like log(7)25x=log(7)100

OpenStudy (anonymous):

its the second equation you typed

Nnesha (nnesha):

yes that's correct i thought 2 represent the number of question like question #2 \[\large\rm \color{blue}{y} \log_b \color{red}{ x}= \log_b \color{Red}{x}^\color{blue}{y}\]exponent rule

Nnesha (nnesha):

\[\log_7 (25x)= \log_7 100\]

OpenStudy (anonymous):

Oh okay thank you for your help though! i got x=4

Nnesha (nnesha):

looks good. good job!